# Flash Diffusion: Accelerating Any Conditional Diffusion Model for Few Steps Image Generation

### v2

@@ 4.2 Timesteps Sampling
where the mean of each Gaussian is controlled by $\{\mu_{i}=i/K\}_{i=1}^{K}$ and the variance is fixed to $\sigma=\sqrt{0.5/K^{2}}$. This approach is such that when distilling the teacher only a small number of $K$ discrete timesteps will be sampled instead of the continuous range $[0,1]$ [^1] In practice when training a DM, the range $[0,1]$ is actually discretized (typically into 1000 timesteps) for computational purposes.. Moreover, the distribution $\pi$ is defined such that out of the $K$ selected timesteps, the 4 timesteps used at inference for 1, 2 and 4 steps generation are over-sampled (typically we set $\beta_{i}>0$ if $i\in[\frac{K}{4},\frac{K}{2},\frac{3K}{4},K]$ and $\beta_{i}=0$ otherwise). Unlike other methods [61, 62] we do not only focus on those 4 timesteps since we noticed that it can lead to a reduction of diversity in the generated samples. This is in particular emphasized in the ablation study presented in Sec. 5.2. In practice, we notice that a warm-up phase is beneficial to the training process. Therefore, we decide to start by first imposing a higher probability to the timesteps corresponding to the least added amount of noise by setting $\beta_{K/4}=\beta_{K/2}=0.5$ and $\beta_{i}=0$ otherwise. We then progressively shift the probability mass towards full noise to favor single-step generation while still over-sampling the targeted 4 timesteps by setting a strictly positive value for $\beta_{i}$ where $i\equiv 0[K/4]$, and $\beta_{i}=0$ otherwise. An example for $\pi$ with $K=32$ is illustrated in Figure 3. As pictured in the figure, the $[0,1]$ interval is split into 32 timesteps. During the warm-up phase, the probability mass allocates a higher probability to timesteps $[0.25,0.5]$ to ease the distillation process. As the training progresses, the probability mass function is then shifted towards full noise to favor single-step generation while always allocating a higher probability to the 4 timesteps $[0.25,0.5,0.75,1]$. The impact of the timesteps distribution is further discussed in Sec. 5.2.
@@
Figure 4: Flash Diffusion training method: the student is trained by using a distillation loss between multiple-step teacher and single-step student denoised samples. The student predictions are then re-noised and denoised with the teacher and student before evaluating the adversarial and distribution matching losses.
4.3 Adversarial Objective
To further enhance the quality of the samples, and since it proved very efficient in several works proposed in the literature to achieve few steps image generation [76, 61, 33, 28, 62], we have also decided to incorporate an adversarial objective. The core idea is to train the student model to generate samples that are indistinguishable from the true data distribution $p(x_{0})$. To do so, we propose to train a discriminator $D_{\nu}$ to distinguish the generated samples $\tilde{x}_{0}$ from the real samples $x_{0}\sim p(x_{0})$. As proposed in [33, 62], we also apply the discriminator directly within the latent space. This approach circumvents the necessity of decoding the samples using the VAE, a process outlined in [61], that proves to be expensive and hampers the method’s scalability to high-resolution images.
@@ 4.3 Adversarial Objective
Drawing inspiration from [33, 62], we propose an approach where both the one-step student prediction $\tilde{z}_{0}$ and the input latent sample $z_{0}$ are re-noised following the teacher noise schedule. This process uses a timestep $t^{\prime}$, which is uniformly chosen from the set $[0.01,0.25,0.5,0.75]$. The samples are first passed through the frozen teacher model, followed by the discriminator, to yield a real or fake prediction. When employing a UNet architecture [58] for the teacher model, our approach focuses on utilizing only the encoder portion of the UNet, generating an even more compressed latent representation and further reducing the parameter count for the discriminator. We carefully select specific timesteps to enable the discriminator to effectively differentiate between samples based on both high and low-frequency details, as discussed by [33]. Note that in our proposed method, the discriminator is the only component we train, while the teacher model remains frozen. The adversarial loss $\mathcal{L}_{\mathrm{adv}}$ and discriminator loss $\mathcal{L}_{\mathrm{discriminator}}$ write as follows:
@@
$\mathcal{L}_{\mathrm{adv}}=$ $~{}\frac{1}{2}~{}\mathbb{E}_{z_{0},t^{\prime},\varepsilon}\left[\left\|D_{\nu}(f_{\theta}^{\mathrm{student}}(z_{t^{\prime}},t^{\prime}))-1\right\|^{2}\right]\,,$ (8)
@@
$\mathcal{L}_{\mathrm{discriminator}}=$ $~{}\frac{1}{2}~{}\mathbb{E}_{z_{0},t^{\prime},\varepsilon}\left[\left\|D_{\nu}(z_{0})-1\right\|^{2}+\left\|D_{\nu}(f_{\theta}^{\mathrm{student}}(z_{t^{\prime}},t^{\prime}))-0\right\|^{2}\right]\,.$
@@
where $\nu$ denotes the discriminator parameters. We opt for these particular losses due to their reliability and stability during training, as observed in our experiments. Additionally, we stress the impact of the chosen adversarial loss $\mathcal{L}_{\mathrm{adv}}$ through an ablation study detailed in Section 5.2. In practical terms, the discriminator’s architecture is designed as a straightforward Convolutional Neural Network (CNN) featuring a stride of 2, a kernel size of 4, SiLU activation [13, 53] and group normalization [75].
4.4 Distribution Matching
@@
Inspired by the work of [77], we also propose to introduce a Distribution Matching Distillation (DMD) loss to ensure that the generated samples closely mirror the data distribution learned by the teacher [12, 9, 31]. Specifically, this involves minimizing the Kullback–Leibler (KL) divergence between the distribution of samples from the student model $p^{\mathrm{student}}_{\theta}$ and $p^{\mathrm{teacher}}_{\phi}$, the data distribution learned by the teacher model [73]:
$\mathcal{L}_{\mathrm{DMD}}$ $=~{}D_{KL}(p^{\mathrm{student}}_{\theta}||p^{\mathrm{teacher}}_{\phi})$ (9)
$=~{}\mathbb{E}_{z_{0},t,\varepsilon}\bigg{[}-\bigg{(}\log p^{\mathrm{teacher}}_{\phi}\big{(}f_{\theta}^{\mathrm{student}}(z_{t},t)\big{)}-\log p^{\mathrm{student}}_{\theta}\big{(}f_{\theta}^{\mathrm{student}}(z_{t},t)\big{)}\bigg{)}\bigg{]}\,.$
@@
Taking the gradient of the KL divergence with respect to the student model parameters $\theta$ leads to the following update rule:
@@
$\nabla_{\theta}\mathcal{L}_{\mathrm{DMD}}=\mathbb{E}_{z_{0},t,\varepsilon}\bigg{[}-\bigg{(}s^{\mathrm{teacher}}\big{(}f_{\theta}^{\mathrm{student}}(z_{t},t)\big{)}-s^{\mathrm{student}}\big{(}f_{\theta}^{\mathrm{student}}(z_{t},t)\big{)}\bigg{)}\nabla f_{\theta}^{\mathrm{student}}(z_{t},t)\bigg{]}\,,$
@@
where $s^{\mathrm{teacher}}$ and $s^{\mathrm{student}}$ are the score functions of the teacher and student distributions respectively. Inspired by [77], the one-step student prediction is re-noised using a uniformly sampled timestep $t^{\prime\prime}\sim\mathcal{U}([0,1])$ and the teacher noise schedule. The new noisy sample is passed through the frozen teacher model to get the score function for the teacher distribution $s^{\mathrm{teacher}}(f_{\theta}^{\mathrm{student}}(z_{t^{\prime\prime}},t^{\prime\prime}))=-(\varepsilon_{\phi}^{\mathrm{teacher}}(x_{t^{\prime\prime}},t^{\prime\prime})/\sigma(t^{\prime\prime}))$. In our approach, we utilize the student model for the score function of the student distribution, instead of a dedicated diffusion model as referenced in [77]. This choice significantly reduces the number of trainable parameters and computational costs. Moreover, the use of the distribution matching loss, examined in the ablation study section (Sec. 5.2), has proven to enhance the quality of generated samples and their prompt adherence.
Algorithm 1 Flash Diffusion
Input: A trained teacher DM $\varepsilon_{\phi}^{\mathrm{teacher}}$, a trainable student DM $\varepsilon_{\theta}^{\mathrm{student}}$, an ODE solver $\Psi$, the number of sampling teacher steps $K$, a timesteps distribution $\pi(t)$, the guidance scale range $[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$, $\lambda_{\mathrm{adv}}$, $\lambda_{\mathrm{dmd}}$ the losses weights
Initialisation: $\theta\leftarrow\phi$$\triangleright$ Initialise the student with teacher’s weights
while not converged do
     $(z,c)\sim\mathcal{Z}\times\mathcal{C}$, $\omega\sim\mathcal{U}\big{(}[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]\big{)}$$\triangleright$ Draw a sample and guidance scale
     $t_{i}\sim\pi(t)$, $\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)$$\triangleright$ Sample a timestep and noise
     $\tilde{z}_{t_{i}}\leftarrow\alpha(t_{i})\cdot z_{0}+\sigma(t_{i})\cdot\varepsilon$
     for $j=i-1\rightarrow 0$ do
         $\tilde{\varepsilon}=\omega\cdot\varepsilon_{\phi}^{\mathrm{teacher}}(\tilde{z}_{t_{j+1}},t_{j+1},c)+(1-\omega)\cdot\varepsilon_{\phi}^{\mathrm{teacher}}(\tilde{z}_{t_{j+1}},t_{j+1},\varnothing)$$\triangleright$ CFG
         $\tilde{z}_{t_{j}}\leftarrow\Psi(\tilde{\varepsilon},t_{j+1},\tilde{z}_{t_{j+1}})$$\triangleright$ ODE solver update
     end for
     $\tilde{z}_{0}^{\mathrm{teacher}}\leftarrow\tilde{z}_{t_{0}}$
     $\tilde{z}_{0}^{\mathrm{student}}\leftarrow\big{(}\tilde{z}_{t_{i}}-\sigma(t_{i})\cdot\varepsilon_{\theta}^{\mathrm{student}}(\tilde{z}_{t_{i}})\big{)}/\alpha(t_{i})$
     $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{distil}}(\tilde{z}_{0}^{\mathrm{student}},\tilde{z}_{0}^{\mathrm{teacher}})$ + $\lambda_{\mathrm{adv}}\cdot\mathcal{L}_{\mathrm{adv}}(\tilde{z}_{0}^{\mathrm{student}},z_{0})$ + $\lambda_{\mathrm{dmd}}\cdot\mathcal{L}_{\mathrm{DMD}}(\tilde{z}_{0}^{\mathrm{student}})$
end while
4.5 Model Training
While striving for robustness and versatility, we also aimed to design a model with a minimal number of trainable parameters, since it involves the loading of computationally intensive functions (teacher and student). To do so, we propose to rely on the parameter-efficient method LoRA [20] and apply it to our student model. This way, we drastically reduce the number of parameters and speed up the training process.
The training process is detailed in Alg. 1 and illustrated in Figure 4. In more detail, we first pick a random sample $x_{0}\sim p(x_{0})$ belonging to the unknown data distribution. This sample is then encoded with an encoder $\mathcal{E}$ to get the corresponding latent sample $z_{0}$. A timestep $t$ is drawn according to the timesteps probability mass function $\pi$ detailed in Sec. 4.2 to create a noisy sample $z_{t}$ using Eq. (1). The teacher model $\varepsilon_{\phi}^{\mathrm{teacher}}$ and the ODE solver $\Psi$ are then used to solve the PF-ODE and so generate a synthetic sample $\tilde{z}_{0}^{\mathrm{teacher}}$ belonging to the distribution learned by the teacher model. At the same time, the student model $f_{\theta}^{\mathrm{student}}$ is used to generate a denoised sample $\tilde{z}_{0}^{\mathrm{student}}=f_{\theta}^{\mathrm{student}}(z_{t},t)$ in a single step. The distillation loss is then computed according to Eq. (6). Then, we re-noise the one-step student prediction $\tilde{z}_{0}^{\mathrm{student}}$ as well as the input latent sample $z_{0}$ and compute the adversarial loss as explained in Sec. 4.3. Finally, for distribution matching, we take again the one-step student prediction $\tilde{z}_{0}^{\mathrm{student}}$ and re-noise it using a uniformly sampled timestep $t\sim\mathcal{U}([0,1])$. The new noisy sample is passed through the teacher model to get the teacher score $s^{\mathrm{teacher}}$ function while we use the student model (and not a dedicated diffusion model as in [77]) to get the student score function $s^{\mathrm{student}}$. The distribution matching loss is then computed as explained in Sec. 4.
Overall, our proposed method relies on the training of only a few number of parameters. This is achieved through applying LoRA to the student model, utilizing a frozen teacher model for the adversarial approach, and employing the student denoiser directly rather than introducing a new diffusion model to calculate the fake scores for the distribution matching loss. This approach not only drastically cuts down on the number of parameters but also accelerates the training process.
5 Experiments
In this section, we assess the effectiveness of our proposed method across various tasks and datasets. First, as it is common in the literature, we quantitatively compare the method with several approaches in the context of text-to-image generation. Then, we conduct an extensive ablation study to assess the importance and impact of each of the components proposed in the method. Finally, we highlight the versatility of our method across several tasks, conditioning, and denoiser architectures.
5.1 Text-to-Image Quantitative Evaluation
First, we propose to evaluate the proposed method against existing distillation approaches for text-to-image generation. In this section, we apply our distillation approach to the publicly available SD1.5 model [57] and report both FID [14] and CLIP score [51] on the COCO2014 and COCO2017 datasets [34]. The model is trained on the LAION dataset [63] where we select samples with aesthetic scores above 6 and re-caption the samples using CogVLM [71]. For COCO2017, we rely on the evaluation approach proposed in [45] and we pick 5,000 prompts from the validation set to generate synthetic images. For COCO2014, we employ the evaluation protocol proposed in [22] and pick 30,000 prompts from the validation set. We then compute the FID against the real images in the respective validation sets composed of 5,000 images for COCO2017 and 40,504 images for COCO2014. The model is trained for only 20k iterations on 2 H100-80Gb GPUs (amounting to a total 26 H100 equivalent hours of training) with a batch size of 4 and a learning rate of $1e^{-5}$ together with Adam optimizer [26] for both the student and the discriminator. We use the timestep distribution $\pi(t)$ detailed in Sec. 4.2 with $K=32$ and shift phases every 5000 iterations. We also start with both $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The guidance scale $\omega$ is sampled from $\mathcal{U}\big{(}[3,13]\big{)}$. The weights of the student model are all initialized with the teacher’s. See Appendix A.1 for more details on the training procedure.
We report the results in Tables 2 and 2. As illustrated in the tables, our method achieves a FID of 22.6 and 12.27 on COCO2017 and COCO2014 respectively with only 2 NFEs corresponding to SOTA results for few steps image generation. On COCO2017, our approach also achieves a CLIP score of 0.306 and 0.311 for 2 and 4 NFEs respectively. Importantly, our method only requires the training of 26.4M parameters (out-of the 900M teacher parameters) and merely 26 H100 GPUs hours of training time. This is in stark contrast with many competitors who depend on training the entire UNet architecture of the student, which involves hundreds of millions of parameters. Beyond quantitative assessments, we also offer a visual overview of the generated samples in Figure 5 for 1, 2, and 4 Network Function Evaluations. In addition to those results, we also provide a quantitative analysis with different backbones (SDXL and Pixart-$\alpha$) in Sec. 5.3.1.
Table 1: FID-5k and CLIP score on MS COCO2017 validation set for SD1.5 as teacher. We follow the procedure of [45] and pick the 5,000 prompt of the validation set.
Method # Train. NFE FID $\downarrow$ CLIP $\uparrow$
Param. (M)
Teacher (SD1.5) N/A 50 20.1 0.318
16 31.7 0.320
Prog. Distil.[45] 900M 2 37.3 0.270
4 26.0 0.300
8 26.9 0.300
InstaFlow [36] 900M 1 23.4 0.304
CFG Distil. [30] 850M 16 24.2 0.300
Ours 26.4 M 2 22.6 0.306
4 22.5 0.311
Table 2: FID-30k on MS COCO2014 validation set for SD1.5 as teacher. We follow the evaluation procedure of [22]. results extracted from [77]
Method # Train. NFE FID $\downarrow$
Param. (M)
DPM++ [38] N/A 8 22.44
UniPC [82] N/A 8 23.30
UFOGen [76] 1,700M 1 12.78
InstaFlow [36] 900M 1 13.10
DMD [77] 1,700M 1 14.93
LCM-LoRA [43] 67.5M 1 77.90
2 24.28
4 23.62
Ours 26.4 M 2 12.27
4 12.41
@@ 5.1 Text-to-Image Quantitative Evaluation
Figure 5: Qualitative evaluation of the sample quality as the number of NFEs increases for the proposed method applied to SD1.5 model. As expected, the quality of the samples increases with the number of NFEs inline with the quantitative results.
5.2 Ablation Study
In the following section, we conduct a comprehensive ablation study to assess the influence of the main parameters and choices made in the proposed method. Part of the ablated parameters are the distillation loss choice, the GAN loss choice, the impact of each term in the loss given in Eq. (10), the timestep sampling as well as the guidance scale used during training. For all the ablations, we train the model for 20k iterations with SD1.5 model as a teacher and employ the same hyper-parameters as in Sec. 5.1 unless stated otherwise. All the results are reported on the COCO2017 validation set using 2 NFEs.
@@ 5.2 Ablation Study
$\mathcal{L}_{\mathrm{adv.}}$ FID $\downarrow$ CLIP $\uparrow$
Hinge 25.02 30.17
WGAN 24.58 30.36
LSGAN 22.64 30.61
$\mathcal{L}_{\mathrm{distil.}}$ FID $\downarrow$ CLIP $\uparrow$
LPIPS 24.89 30.56
MSE 22.64 30.61
$K$ FID $\downarrow$ CLIP $\uparrow$
16 23.35 30.52
32 22.64 30.61
64 22.87 30.41
@@
Figure 6: From left to right and top to bottom: Influence of the loss terms, the GAN loss choice, the distillation loss choice, the guidance scale used to generate with the teacher, $K$ the number of reference timesteps and the timestep sampling $\pi(t)$,.
@@
First, we assess the influence of each of the loss terms in the method’s loss given in Eq. (10). In this ablation study, we train the model using various combinations of loss terms: solely the distillation loss, the distillation loss combined with the DMD loss, the distillation loss paired with the adversarial loss, or incorporating all loss terms simultaneously. We report the results in the top left table in Figure 6. As highlighted in the table, the distillation loss alone leads to higher FID and lower CLIP score compared to the other configurations showcasing the importance of the adversarial and DMD losses. Moreover, both losses have a noticeable impact on the final performance since the adversarial loss seems to allow reaching a better image quality, as indicated by lower FID, while the DMD loss improves prompt adherence, reflected in higher CLIP scores. Experiments conducted using only adversarial and DMD losses revealed notable inconsistencies and even divergence in outcomes, emphasizing the crucial contribution of the distillation loss to the method’s stability and reliability. This highlights the importance of a balanced approach that incorporates all three losses for optimal results.
In the top middle and top right tables in Figure 6, we also report some results obtained with a model trained with different choices for the distillation loss and adversarial loss. For the distillation loss, we compare the use of LPIPS [81] and MSE. As highlighted in the table, the use of MSE allows to achieve better results in terms of FID and CLIP score than the use of LPIPS. For the GAN loss, we decide to compare the use of a Hinge loss [32], the approach of WGAN [1] or LSGAN [44]. As illustrated in the table, the use of LSGAN seems the best-suited choice. Moreover, we noticed stabler trainings using this loss than counterparts.
@@
As explained in section 4.2, we select $K=32$ uniformly spaced timesteps in the range $[0,1]$ and assign a probability to each of them according to a probability mass function given by $\pi(t)$. In this section, we highlight the influence of the choice for $\pi(t)$. We compare the proposed timestep distribution described in Sec. 4.2 to the use of a uniform distribution across those 32 timesteps $\pi^{\mathrm{uniform}}(t)=\mathcal{U}\left(\{i/K\}_{i\in\{1,\dots,K\}}\right)$, a normal distribution $\pi^{\mathrm{gaussian}}(t)$ centered on $t=0.5$ and $\pi^{\mathrm{sharp}}$, a sharp version of our proposed distribution that only allows sampling 4 distinct timesteps. We report the results in the bottom right table in Figure 6. These distributions are represented in the Appendix A.1. The first outcome of such an experiment is that, as expected, the choice in the timestep sampling distribution has a noticeable impact on the final performance. As shown by the two bottom lines, the proposed distribution significantly improves the performance compared to the uniform and Gaussian distribution. Moreover, allowing to sample more than 4 distinct timesteps seems to be beneficial to the final performance since a noticeable decrease in the FID score is observed. This can be explained by the fact that the student model can distill more useful information from the teacher model by sampling a wider range of timesteps and not over-fit on the 4 selected ones.
In addition, we also assess the influence of the number of selected timesteps $K$ used to create the timestep distribution $\pi(t)$ described in Figure 3. We compare 3 cases: $K=16$, $K=32$ and $K=64$ and report the results in the right-most table in Figure 6. In this experiment, $K=32$ reveales to be a good trade-off between performance and computational cost. Indeed, the use of $K=16$ seems to lead to a slight decrease in the final performance while $K=64$ also involves a higher computational cost. It indeed induces a doubled number of NFEs used to produce a denoised sample with the teacher during training while the performances are roughly similar to $K=32$.
Influence of the guidance scale

### v3

@@ Introduction
where the mean of each Gaussian is controlled by $\{\mu_{i}=i/K\}_{i=1}^{K}$ and the variance is fixed to $\sigma=\sqrt{0.5/K^{2}}$. This approach is such that when distilling the teacher only a small number of $K$ discrete timesteps will be sampled instead of the continuous range $[0,1]$ [^1] In practice when training a DM, the range $[0,1]$ is actually discretized (typically into 1000 timesteps) for computational purposes.. Moreover, the distribution $\pi$ is defined such that out of the $K$ selected timesteps, the 4 timesteps used at inference for 1, 2 and 4 steps generation are over-sampled (typically we set $\beta_{i}>0$ if $i\in[\frac{K}{4},\frac{K}{2},\frac{3K}{4},K]$ and $\beta_{i}=0$ otherwise). Unlike other methods (Sauer et al. 2023, 2024) we do not only focus on those 4 timesteps since we noticed that it can lead to a reduction of diversity in the generated samples. This is in particular emphasized in the ablation study. In practice, we notice that a warm-up phase is beneficial to the training process. Therefore, we decide to start by first imposing a higher probability to the timesteps corresponding to the least added amount of noise by setting $\beta_{K/4}=\beta_{K/2}=0.5$ and $\beta_{i}=0$ otherwise. We then progressively shift the probability mass towards full noise to favor single-step generation while still over-sampling the targeted 4 timesteps by setting a strictly positive value for $\beta_{i}$ where $i\equiv 0[K/4]$, and $\beta_{i}=0$ otherwise. An example for $\pi$ with $K=32$ is illustrated in Fig. 2. As pictured in the figure, the $[0,1]$ interval is split into 32 timesteps. During the warm-up phase, the probability mass allocates a higher probability to timesteps $[0.25,0.5]$ to ease the distillation process. As the training progresses, the probability mass function is then shifted towards full noise to favor single-step generation while always allocating a higher probability to the 4 timesteps $[0.25,0.5,0.75,1]$. The impact of the timesteps distribution is further discussed in the ablations.
@@
Figure 3: Flash Diffusion training method: the student is trained with a distillation loss between multiple-step teacher and single-step student denoised samples. The student predictions are then re-noised and denoised with the teacher and student before evaluating the GAN and DMD losses.
Adversarial Objective
@@
To further enhance the quality of the samples, we have also decided to incorporate an adversarial objective. The core idea is to train the student model to generate samples that are indistinguishable from the true data distribution $p(x_{0})$. To do so, we propose to train a discriminator $D_{\nu}$ to distinguish the generated samples $\tilde{x}_{0}$ from the real samples $x_{0}\sim p(x_{0})$. As proposed in (Lin, Wang, and Yang 2024; Sauer et al. 2024), we also apply the discriminator directly within the latent space. This approach circumvents the necessity of decoding the samples using the VAE, a process outlined in (Sauer et al. 2023), that proves to be expensive and hampers the method’s scalability to high-resolution images. Drawing inspiration from (Lin, Wang, and Yang 2024; Sauer et al. 2024), we propose an approach where both the one-step student prediction $\tilde{z}_{0}$ and the input latent sample $z_{0}$ are re-noised following the teacher noise schedule. This process uses a timestep $t^{\prime}$ uniformly chosen from the set $[0.01,0.25,0.5,0.75]$ enabling the discriminator to effectively differentiate between samples based on both high and low-frequency details (Lin, Wang, and Yang 2024). The samples are first passed through the frozen teacher model, followed by the trainable discriminator, to yield a real or fake prediction. When employing a UNet architecture (Ronneberger, Fischer, and Brox 2015) for the teacher model, our approach focuses on utilising only the encoder of the UNet, generating an even more compressed latent representation and further reducing the parameter count for the discriminator. The adversarial loss $\mathcal{L}_{\mathrm{adv}}$ and discriminator loss $\mathcal{L}_{\mathrm{dis}}$ write as follows:
@@
$\mathcal{L}_{\mathrm{adv}}=$ $~{}\frac{1}{2}~{}\mathbb{E}_{z_{0},t^{\prime},\varepsilon}\left[\left\|D_{\nu}(f_{\theta}(z_{t^{\prime}},t^{\prime}))-1\right\|^{2}\right]\,,$ (4)
@@
$\mathcal{L}_{\mathrm{dis.}}=$ $~{}\frac{1}{2}~{}\mathbb{E}_{z_{0},t^{\prime},\varepsilon}\left[\left\|D_{\nu}(z_{0})-1\right\|^{2}+\left\|D_{\nu}(f_{\theta}(z_{t^{\prime}},t^{\prime}))\right\|^{2}\right]\,,$
@@
where $\nu$ denotes the discriminator parameters. We opt for these particular losses due to their reliability and stability during training, as observed in our experiments. In practical terms, the discriminator’s architecture is designed as a straightforward Convolutional Neural Network (CNN) featuring a stride of 2, a kernel size of 4, SiLU activation (Hendrycks and Gimpel 2016; Ramachandran, Zoph, and Le 2017) and group normalization (Wu and He 2018).
Distribution Matching
@@
Inspired by the work of (Yin et al. 2023), we also propose to introduce a Distribution Matching Distillation (DMD) loss to ensure that the generated samples closely mirror the data distribution learned by the teacher. Specifically, this involves minimizing the Kullback–Leibler (KL) divergence between the student distribution $p^{\mathrm{student}}_{\theta}$ and $p^{\mathrm{teacher}}_{\phi}$, the data distribution learned by the teacher (Wang et al. 2024):
$\mathcal{L}_{\mathrm{DMD}}$ $=~{}D_{KL}(p^{\mathrm{student}}_{\theta}||p^{\mathrm{teacher}}_{\phi})\,.$ (5)
@@
Taking the gradient of the KL divergence with respect to the student model parameters $\theta$ leads to the following update:
@@
$\nabla_{\theta}\mathcal{L}_{\mathrm{DMD}}=\mathbb{E}\bigg{[}\bigg{(}s^{\mathrm{student}}(y)-s^{\mathrm{teacher}}(y))\big{)}\bigg{)}\nabla f_{\theta}(z_{t},t)\bigg{]}\,,$
@@
where $s^{\mathrm{teacher}}$ and $s^{\mathrm{student}}$ are the score functions of the teacher and student distributions respectively and $y=f_{\theta}(z_{t},t)$ is the student prediction. Inspired by (Yin et al. 2023), the one-step student prediction $\tilde{z}_{0}$ is re-noised using a uniformly sampled timestep $t^{\prime\prime}\sim\mathcal{U}([0,1])$ and the teacher noise schedule. The new noisy sample is passed through the frozen teacher model to get the score function for the teacher distribution $s^{\mathrm{teacher}}(f_{\theta}(z_{t^{\prime\prime}},t^{\prime\prime}))=-(\varepsilon_{\phi}^{\mathrm{teacher}}(x_{t^{\prime\prime}},t^{\prime\prime})/\sigma(t^{\prime\prime}))$. In our approach, we utilize the student model for the score function of the student distribution, instead of a dedicated diffusion model. This choice significantly reduces the number of trainable parameters and computational costs.
@@
Figure 4: Qualitative evaluation of the sample quality as the number of NFEs increases for the proposed method applied to SD1.5 model. Best viewed zoomed in.
Model Training
While striving for robustness and versatility, we also aimed to design a model with a minimal number of trainable parameters, since it involves the loading of computationally intensive functions (teacher and student). To do so, we propose to rely on the parameter-efficient method LoRA (Hu et al. 2021) and apply it to our student model. This way, we drastically reduce the number of parameters and speed up the training process.
The training process is illustrated in Fig. 3 and detailed in the appendices.
Experiments
In this section, we assess the effectiveness of our proposed method across various tasks and datasets. First, as it is common in the literature, we quantitatively compare the method with several approaches in the context of text-to-image generation. Then, we conduct an extensive ablation study to assess the importance and impact of each component proposed in the method. Finally, we highlight the versatility of our method across several tasks, conditioning, and denoiser architectures.
Text-to-Image Quantitative Evaluation
First, we apply our distillation approach to the publicly available SD1.5 model (Rombach et al. 2022) and report both FID (Heusel et al. 2017) and CLIP score (Radford et al. 2021) on the COCO2014 and COCO2017 datasets (Lin et al. 2014). The model is trained on the LAION dataset (Schuhmann et al. 2022) with aesthetic scores above 6. For COCO2017, we rely on the evaluation approach proposed in (Meng et al. 2023) and we pick 5,000 prompts from the validation set to generate synthetic images. For COCO2014, we follow (Kang et al. 2023) and pick 30,000 prompts from the validation set. We then compute the FID against the real images in the respective validation sets. We report the results in Tables (a) and (b) in Fig. 5. Our method achieves a FID of 22.6 and 12.27 on COCO2017 and COCO2014 respectively with only 2 NFEs corresponding to SOTA results for few steps image generation. On COCO2017, our approach also achieves a CLIP score of 30.6 and 31.1 for 2 and 4 NFEs respectively. Importantly, our method only requires the training of 26.4M parameters (out-of the 900M teacher parameters) and merely 26 H100 GPUs hours of training time. This is in stark contrast with many competitors who depend on training the entire UNet architecture of the student. See the appendices for more details on the training procedure.
Ablation Study
In this section, we conduct a comprehensive ablation study to assess the influence of the main parameters and choices made in the proposed method. For all the ablations, we train the model for 20k iterations with SD1.5 model as a teacher. All the results are reported on the COCO2017 using 2 NFEs.
Method (# NFE) # Train. FID $\downarrow$ CLIP $\uparrow$
Param.
SD1.5 (50) N/A 20.1 31.8
SD1.5 (16) 31.7 32.0
Prog. Distil. (2) 900M 37.3 27.0
Prog. Distil. (4) 26.0 30.0
Prog. Distil. (8) 26.9 30.0
InstaFlow (1) 900M 23.4 30.4
CFG Dist. (16) 850M 24.2 30.0
Ours (2) 26.4 M 22.6 30.6
Ours (4) 22.5 31.1
Method (# NFE) # Train. FID $\downarrow$
Param.
DPM++ (8) N/A 22.44
UniPC (8) N/A 23.30
UFOGen (1) 1,700M 12.78
InstaFlow (1) 900M 13.10
DMD (1) 1,700M 14.93
LCM-LoRA (1) 67.5M 77.90
LCM-LoRA (2) 24.28
LCM-LoRA (4) 23.62
Ours (2) 26.4M 12.27
Ours (4) 12.41
@@
$\mathcal{L}_{\mathrm{distil.}}$ FID $\downarrow$ CLIP $\uparrow$
LPIPS 24.89 30.56
MSE 22.64 30.61
$\mathcal{L}_{\mathrm{adv.}}$ FID $\downarrow$ CLIP $\uparrow$
Hinge 25.02 30.17
WGAN 24.58 30.36
LSGAN 22.64 30.61
$K$ FID $\downarrow$ CLIP $\uparrow$
16 23.35 30.11
32 22.64 30.61
64 22.87 30.58
Figure 5: From left to right and top to bottom: a) FID-5k and CLIP score on COCO2017 validation set for SD1.5 as teacher. b) FID-30k on MS COCO2014 validation set for SD1.5 as teacher ( results from (Yin et al. 2023)). c) Influence of the guidance scale used to generate with the teacher, d) the loss terms e) the timestep sampling $\pi(t)$, f) the distillation loss, g) the GAN loss and h) the value of $K$ in Eq. (3).
@@
We first train the model using different loss combinations and report the results in Table (d) in Figure 5. As highlighted in the table, both $\mathcal{L}_{\mathrm{adv}}$ and $\mathcal{L}_{\mathrm{DMD}}$ have a noticeable impact on the final performance since $\mathcal{L}_{\mathrm{adv}}$ seems to allow reaching a better image quality, as indicated by lower FID, while $\mathcal{L}_{\mathrm{DMD}}$ improves prompt adherence, reflected in higher CLIP scores. Experiments conducted using only $\mathcal{L}_{\mathrm{adv}}$ and $\mathcal{L}_{\mathrm{DMD}}$ revealed notable inconsistencies and even divergence in outcomes, emphasizing the crucial contribution of the distillation loss to the method’s stability and reliability. In Tables (f) and (g), we also report results for different $\mathcal{L}_{\mathrm{distil.}}$ (LPIPS (Zhang et al. 2018) and MSE) and $\mathcal{L}_{\mathrm{adv}}$ (Hinge (Lim and Ye 2017), WGAN (Arjovsky, Chintala, and Bottou 2017) and LSGAN (Mao et al. 2017)). For $\mathcal{L}_{\mathrm{distil.}}$, MSE allows to achieve better results in terms of FID and CLIP score than LPIPS. For the GAN loss, the use of LSGAN seems the best-suited choice and we also noticed that it leads to stabler trainings.
@@
In this section, we stress the influence of $\pi(t)$, the timesteps distribution. We compare the proposed timestep distribution to a uniform distribution across $K=32$ timesteps, a normal distribution $\pi^{\mathrm{gaussian}}(t)$ centered on $t=0.5$ and $\pi^{\mathrm{sharp}}$, a sharp version of our proposed distribution that only allows sampling 4 distinct timesteps. Results are shown in Table (e) of Fig. 5. The proposed distribution significantly improves the performance compared to $\pi^{\mathrm{uniform}}$ and $\pi^{\mathrm{gaussian}}$. Moreover, allowing to sample more than 4 distinct timesteps seems to be beneficial to the final performance since a noticeable decrease in the FID score is observed. This can be explained by the fact that the student model can distil more useful information from the teacher model by sampling a wider range of timesteps and not over-fit the 4 selected ones.
Influence of the guidance scale during training